Engine failure forced Robbin to Pilot be Cessna 150 to an emergency landing. To land, Robbins plane glided 17,000 ft over a 15,000-ft stretch of deserted highway. From what altitude did the decent begin?

Thank you

Answer by checkley77(12844)

You have a right triangle with an hypotenuse=17,000, a side=15,000 and an unknown height (the third side of the triangle).
17,000^2=15,000^2+x^2
289,000,000=225,000,000+x^2
x^2=289,000,000-225,000,000
x^2=64,000,000
x=sqrt64,000,000
x=8,000 ft was the altitude when the pilot began his glide.